# FLRONet: Deep Operator Learning for High-Fidelity Fluid Flow Field Reconstruction from Sparse Sensor Measurements

### v3

@@ I Introduction
The ability to reconstruct the fluid flow field from limited sensor measurements is of paramount importance in numerous scientific and engineering applications. In this problem, the objective is to reconstruct the high-fidelity flow field $\mathbf{u}(t)$ from a sparse set of sensor observations $\mathbf{y}(t)$. Here, both $\mathbf{u}(t)$ and $\mathbf{y}(t)$ are functions of time and return values in $\mathbb{R}^{m}$ and $\mathbb{R}^{p}$ (with $m>>p$), respectively. This task is typically difficult due to the distinctive characteristics of the measurement operator $\mathcal{H}:\mathbb{R}^{m}\to\mathbb{R}^{p}$ that maps a specific state of the flow field $\mathbf{u}$ at time $t=\tau$ to its corresponding observations $\mathbf{y}$. Due to the large discrepancy between the dimension of $\mathbb{R}^{m}$, i.e., the state space, and the dimension of $\mathbb{R}^{p}$, i.e., the observational space, the measurement operator $\mathcal{H}$ is often ill conditioned and thus is noninvertible. Therefore, it is nearly impossible to derive the forward map $\mathcal{G}:\mathbb{R}^{p}\to\mathbb{R}^{m}$ as the inverse of $\mathcal{H}$. For this reason, alternative methodologies such as data-driven approaches Erichson et al. (2020), sparse optimization Loiseau, Noack, and Brunton (2018), and data assimilation Mons et al. (2016) are frequently adopted. However, despite significant progress reported in the literature, the reconstruction of fluid flow fields from sparse sensor measurements remains a formidable and unresolved challenge.
@@ II Problem formulation
Figure 1: Problem formulation of FLRONet. FLRONet utilize deep operator network to reconstruct the original flow field evolution for a given observation time window $\left[\tau_{0},\tau_{n}\right]$. The reconstruction is continuous in both spatial and temporal domains.
@@ III.2 Operator network for spatiotemporal discretization independent
$\text{FNO}(\Omega(x))=\mathcal{F}^{-1}(\mathcal{F}(\kappa)\cdot\mathcal{F}(\Omega))(x),$ (3)
@@
Considering that we are restricted to evaluating the value $\mathbf{u}(x)$ on a finite mesh. Equations 4 can be discretized by the two-dimensional discrete Fourier transform (DFT) and the inverse discrete Fourier transform (IDFT). Assume that the examined domain is discretized into a uniform grid, where $N_{h}$ and $N_{w}$ are the number of grid points in the vertical and horizontal dimensions, respectively. The two-dimensional DFT and its inverse IDFT for a function $f$ defined on this grid are given by:
@@ III.3 Voronoi tessellation for sensor measurements embedding
Traditional regressive field reconstruction methods often represent sensor measurements as vectors Erichson et al. (2020); Dubois et al. (2022), restricting the ability of machine learning models to generalize effectively to different sensor settings (e.g., varying sensor locations or layouts). Particularly, this representation scheme omits information about sensor positions, leaving ML models unaware of sensor locations and causing them to struggle to adapt to different sensor configurations, e.g., different sensor position Cheng et al. (2024). Several pioneer studies have attempted to apply masked images as a way to encode sensor positions; however, these approaches cannot achieve proper reconstruction accuracy, especially when sensor measurements are sparse. In this work, we used the Voronoi tessellation to encode the sensor measurement information.
@@ III.4 Data and Training
We trained FLRONet using the CFDBench cylinder dataset Yining, Yingfa, and Zhen (2023). The dataset contains numerical simulations of fluid flow around a cylindrical obstacle (Fig. 4). The examined domain has a physical size of $0.14\times 0.24(m)$, which is tessellated in a grid of $140\times 240$ (pixels). There are 50 simulation cases in the dataset, each having a unique inlet velocity increasing incrementally from 0.1 m/s to 5.0 m/s. For each simulation case, the total simulation time of 1 second is recorded using 1,000 snapshots with the time interval between snapshots being 0.001 (s). For each case, 32 sensors are randomly distributed across the domain to record velocity magnitudes at 32 different locations, providing sparse observations that FLRONet uses as input. Fig. 4 illustrates the dimensions of the examined domain and the sensor layout.
@@
To construct the training data, we created observation windows containing five snapshots of sensor data, with the time interval between two snapshots being 0.005 (s), resulting in a window of 0.02-seconds. With this setup, 980 observation windows were generated for a single simulation case. Within each window, we randomly select a single snapshot to record the velocity field at that time point, which serves as the ground truth for FLRONet. To evaluate generalization, we randomly sampled the 50 simulation cases into two exclusive sets: 45 cases for training and five cases for testing. The training set includes all cases of inlet velocity, except for those specifically reserved for the test set. The test set consists of cases with inlet velocities of 3.5, 3.9, 4.2, 4.6, and 5.0 m/s. FLRONet was trained using the Adam optimizer with a learning rate of $10^{-3}$.
@@ IV.4 Zero-shot super resolution in spatial domain
Figure 9: Super-resolution in space> The top row shows the full-field reconstruction, the bottom row provides magnified views of the highlighted regions to illustrate the improvement in spatial detail. Here, we can see the progressive reduction in pixelation and enhancement of spatial details as the resolution increases. Theocratically, FLRONet-FNO is able to reconstruct the complete vector field at arbitrary scales, constrained only by hardware memory.

### v4

@@ I Introduction
The ability to reconstruct the fluid flow field from limited sensor measurements is of paramount importance in numerous scientific and engineering applications. In this problem, the objective is to reconstruct the high-fidelity flow field $\mathbf{u}(t)$ from a sparse set of sensor observations $\mathbf{y}(t)$. Here, both $\mathbf{u}(t)$ and $\mathbf{y}(t)$ are functions of time and return values in $\mathbb{R}^{m}$ and $\mathbb{R}^{p}$ (with $m\gg p$), respectively. This task is typically difficult due to the distinctive characteristics of the measurement operator $\mathcal{H}:\mathbb{R}^{m}\to\mathbb{R}^{p}$ that maps a specific state of the flow field $\mathbf{u}$ at time $t=\tau$ to its corresponding observations $\mathbf{y}$. Due to the large discrepancy between the dimension of $\mathbb{R}^{m}$, i.e., the state space, and the dimension of $\mathbb{R}^{p}$, i.e., the observational space, the measurement operator $\mathcal{H}$ is often ill conditioned and thus is noninvertible. Therefore, it is nearly impossible to derive the forward map $\mathcal{G}:\mathbb{R}^{p}\to\mathbb{R}^{m}$ as the inverse of $\mathcal{H}$. For this reason, alternative methodologies such as data-driven approaches Erichson et al. (2020), sparse optimization Loiseau, Noack, and Brunton (2018), and data assimilation Mons et al. (2016) are frequently adopted. However, despite significant progress reported in the literature, the reconstruction of fluid flow fields from sparse sensor measurements remains a formidable and unresolved challenge.
@@ II Problem formulation
Figure 1: Problem formulation of FLRONet. FLRONet utilizes deep operator network to reconstruct the original flow field evolution for a given observation time window $\left[\tau_{0},\tau_{n}\right]$. The reconstruction is continuous in both spatial and temporal domains.
@@ III.2 Operator network for spatiotemporal discretization independent
$\text{FNO}(\Omega(\mathbf{x}))=\mathcal{F}^{-1}(\mathcal{F}(\kappa)\cdot\mathcal{F}(\Omega))(\mathbf{x}),$ (3)
@@
Considering that we are restricted to evaluating the value $\mathbf{u}(\mathbf{x})$ on a finite mesh. Equations 4 can be discretized by the two-dimensional discrete Fourier transform (DFT) and the inverse discrete Fourier transform (IDFT). Assume that the examined domain is discretized into a uniform grid, where $N_{h}$ and $N_{w}$ are the number of grid points in the vertical and horizontal dimensions, respectively. The two-dimensional DFT and its inverse IDFT for a function $f$ defined on this grid are given by:
@@ III.3 Voronoi tessellation for sensor measurements embedding
Traditional regressive field reconstruction methods often represent sensor measurements as vectors Erichson et al. (2020); Dubois et al. (2022), restricting the ability of machine learning models to generalize effectively to different sensor settings (e.g., varying sensor locations or layouts). Particularly, this representation scheme omits information about sensor positions, leaving ML models unaware of sensor locations and causing them to struggle with adapting to different sensor configurations, e.g., different sensor positions Cheng et al. (2024). Several pioneer studies attempted to apply masked images as a way to encode sensor positions; however, these approaches cannot achieve proper reconstruction accuracy, especially when sensor measurements are sparse. In this work, we used the Voronoi tessellation to encode the sensor measurement information.
@@ III.4 Data and Training
We trained FLRONet using the CFDBench cylinder dataset Yining, Yingfa, and Zhen (2023). The dataset contains numerical simulations of fluid flow around a cylindrical obstacle (Fig. 4). The examined domain has a physical size of $0.14\times 0.24(m)$, which can be tessellated in a grid of $140\times 240$ (pixels). There are 50 simulation cases in the dataset, each having a unique inlet velocity increasing incrementally from 0.1 m/s to 5.0 m/s. For each simulation case, the total simulation time is 1 second, which is recorded using 1,000 snapshots. The time interval between snapshots is 0.001 seconds. For each case, 32 sensors are randomly distributed across the domain to record velocity magnitudes at 32 different locations, providing sparse observations that FLRONet uses as input. Fig. 4 illustrates the dimensions of the examined domain and the sensor layout.
@@
To construct the training data, we created observation windows containing five snapshots of sensor data, with the time interval between two snapshots being 0.005 seconds, resulting in a window of 0.02 seconds. With this setup, 980 observation windows were generated for a single simulation case. Within each window, we randomly select a single snapshot to record the velocity field at that time point, which serves as the ground truth for FLRONet. To evaluate generalization, we randomly sampled the 50 simulation cases into two exclusive sets: 45 cases for training and five cases for testing. The training set includes all cases of inlet velocity, except for those specifically reserved for the test set. The test set consists of cases with inlet velocities of 3.5, 3.9, 4.2, 4.6, and 5.0 m/s. FLRONet was trained using the Adam optimizer with a learning rate of $10^{-3}$.
@@ IV.4 Zero-shot super resolution in spatial domain
Figure 9: Super-resolution in space The top row shows the full-field reconstruction, the bottom row provides magnified views of the highlighted regions to illustrate the improvement in spatial detail. Here, we can see the progressive reduction in pixelation and enhancement of spatial details as the resolution increases. Theocratically, FLRONet-FNO is able to reconstruct the complete vector field at arbitrary scales, constrained only by hardware memory.